OpenStudy (username123):

SOMEBODY ANYBODY EVERYBODY HELP!! Identify the correct slope and y intercept of the equation x - 4y = 4.

OpenStudy (strawberry17):

x - 4y = 4 -x -x ---------- -4y = -x + 4 -4y = -x + 4 --- ----- -4 -4 1 y = --x - 1 4 ^^This is the y intercept. The slope is -1.

OpenStudy (precal):

slope is 1/4 and your y intercept is -1 everything else is correct

OpenStudy (strawberry17):

that't true, my mistake. Thanks

OpenStudy (precal):

 yes not sure if your two dashes are the division bar, almost thought it represented a negative fraction